# Evaluating Human-AI Collaboration: A Review and Methodological Framework

### v1

@@ Abstract
The use of artificial intelligence (AI) in working environments with individuals, known as Human-AI Collaboration (HAIC), has become essential in a variety of domains, boosting decision-making, efficiency, and innovation. Despite HAIC’s wide potential, evaluating its effectiveness remains challenging due to the complex interaction of components involved. This paper provides a detailed analysis of existing HAIC evaluation approaches and develops a fresh paradigm for more effectively evaluating these systems. Our framework includes a structured decision tree which assists to select relevant metrics based on distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic). By including both quantitative and qualitative metrics, the framework seeks to represent HAIC’s dynamic and reciprocal nature, enabling the assessment of its impact and success. This framework’s practicality can be examined by its application in an array of domains, including manufacturing, healthcare, finance, and education, each of which has unique challenges and requirements. Our hope is that this study will facilitate further research on the systematic evaluation of HAIC in real-world applications.
Keywords Human-AI Collaboration (HAIC)  $\cdot$ HAIC Evaluation  $\cdot$ Symbiotic AI Systems  $\cdot$ AI Metrics  $\cdot$ AI Evaluation Framework
@@ 1 Introduction
The fast advancement of Artificial Intelligence (AI) has fundamentally transformed several domains, blurring the distinctions between human and machine roles and giving rise to new interactions [1]. The notion of Human-AI Collaboration (HAIC) has emerged as a pivotal framework, as evidenced by the research conducted by [2] and [3]. These studies identify that HAIC has significant potential for enhancing decision-making, increasing efficiency, and fostering innovation. They therefore aim at fostering and facilitating effective cooperation between people and AI systems to achieve common goals. On the other hand, assessing the benefits of this cooperation presents its own challenges, owing to the multiple complex factors involved.
Traditional Human-Machine evaluation frameworks have primarily focused on quantifiable metrics such as task performance, response time, and user satisfaction within human-machine systems [4, 5]. These frameworks typically view the machine as a tool and assess how effectively humans can control or interact with this tool to achieve desired outcomes.
Human-AI Collaboration (HAIC) represents a significant change in which humans and AI work together in a more integrated and mutually beneficial manner. HAIC focuses on more than just human dominance over machines; it emphasizes a collaborative partnership where individuals and AI systems actively contribute to shared results. The significance of this partnership has been highlighted in several domains, such as manufacturing [6, 7], financial services [8, 9], healthcare [10, 11], education [12], and others, demonstrating notable advancements in both technological innovation and practical implementation [13, 14].
HAIC is defined by the process of making decisions together, learning from each other, and adapting, which requires a more comprehensive and nuanced evaluation approach [15, 16]. To effectively evaluate HAIC, it is crucial to consider not only the traditional metrics of task success but also the quality of interactions within the team, the processes by which decisions are made, and crucial ethical aspects such as the transparency of algorithms and the mitigation of biases [17]. This involves assessing how AI systems influence human decisions, how humans and AI adapt to each other’s capabilities, and how ethical considerations are integrated into collaborative processes. In this paper, we aim at addressing these evolving needs by critically analyzing current evaluation methods in human-AI interaction and proposing a new framework designed to capture the dynamic and reciprocal nature of human-AI collaborations. This framework seeks to recognize and measure both the individual and joint contributions of humans and AI to collaborative outcomes, thus providing a more comprehensive understanding of HAIC effectiveness.
The contributions of this paper are therefore the following.
  • Overview of the Literature: Extending the foundational work of Madni [2] and Dubey et al. [3], we provide a review of HAIC evaluation methods, identifying gaps and research opportunities in the existing literature.
  • Novel Evaluation Framework: Inspired by the pioneering frameworks of Rouse et al.[18] and Parasuraman et al.[19, 4], we propose a new, adaptable evaluation framework for HAIC that incorporates both quantitative and qualitative measures, emphasizing ethical considerations and providing a flexible structure for diverse HAIC applications.
  • Domain-Specific Insights: Through detailed examination of applications in healthcare, finance, and education, as informed by the studies of Tschandl et al.[10], Buckley and Lui[8, 9], and Holstein et al. [12], we demonstrate the applicability and customization of our framework to meet domain-specific needs.
Structure of the Paper
Following this introduction, Section 2 discusses the fundamentals of Human-AI Collaboration, including various modes of interaction. Section 3 reviews the existing literature on HAIC, highlighting advancements in evaluation frameworks and tools, as well as the diversity and challenges in domain-specific evaluations. Section 4 presents a structured framework for evaluating HAIC, detailing factors, subfactors, and metrics for assessment. Section 5 applies this evaluation framework to real-world domains such as manufacturing, healthcare, finance, and education. Section 6 addresses the unique challenges of evaluating creative and linguistic AI, including large language models and generative AI in the arts. Section 7 discusses the limitations in investigating behavioral factors. Finally, Section 8 discusses the implications of the HAIC evaluation framework for future research and practice.
2 Understanding the Fundamentals of Human-AI Collaboration
@@ 2.1 From Human-Machine Interaction to Human-AI Collaboration
Given the rapid pace of advancements in Human-AI Collaboration (HAIC), a comprehensive understanding of the current landscape is essential. While foundational research in Human-Machine Interaction (HMI) provides valuable insights into human-technology interactions, it often falls short when applied to HAIC evaluations. Traditional HMI methods, as exemplified in the works of Nielsen [20], Shneiderman [21], and Norman [22], primarily focus on assessing the usability and effectiveness of tools or interfaces designed to augment human performance. These methods often rely on metrics such as task completion time, error rates, and user satisfaction to gauge the success of human-machine interactions. However, these metrics are not always sufficient to capture the complexities of HAIC, where the relationship between humans and AI is more dynamic and collaborative. For example, traditional metrics may not adequately assess the level of trust a human places in an AI system’s recommendations or the effectiveness of shared decision-making processes where both the human and AI contribute to the final outcome [23]. Additionally, these metrics may not fully capture the adaptability of AI systems, which is crucial for HAIC as the AI needs to learn and adjust its behavior based on human input and changing circumstances [24]. To effectively evaluate HAIC, we must move beyond the limitations of traditional HMI approaches. We need to consider not just the performance of the individual components (human and AI) but also the quality of their collaboration, the fluidity of their interactions, and the achievement of shared goals. This necessitates a broader range of metrics and methodologies that can capture the nuances of these complex partnerships.
2.2 Elements of HAIC
Human-AI Collaboration (HAIC) is the cooperative partnership between individuals and artificial intelligence (AI) systems in order to accomplish common goals.
Several key elements define the landscape of HAIC:
@@ 2.2 Elements of HAIC
  • Tasks: HAIC systems can be designed to tackle a wide array of tasks, from complex decision-making to creative endeavors and knowledge work. The nature of these tasks often dictates the level and type of collaboration required between humans and AI [25, 26]. For example, in the medical field, an AI system can analyze medical images to identify potential anomalies, while a radiologist interprets the results and makes a final diagnosis.
  • Goals: HAIC is driven by shared goals, which can be both individual and collective. Individual goals might encompass AI objectives like improving efficiency and accuracy or human aims like skill enhancement and knowledge acquisition. Collective goals focus on achieving overarching objectives that benefit from the combined capabilities of both humans and AI [27]. For example, in a healthcare setting, an AI system could be designed to examine medical images and identify initial indications of illnesses like cancer (individual goal). The human radiologist review’s the the AI’s findings considers patient history, and makes the final diagnosis (individual goal). The collective goal in this scenario is to improve diagnostic accuracy and ensure timely and effective patient care, thereby enhancing overall health outcomes.
  • Interaction: The success of HAIC heavily relies on effective communication and feedback mechanisms between humans and AI. The quality of this interaction determines how well each party understands the other’s intentions, capabilities, and limitations [17, 28]. For instance, an AI chatbot for customer service needs to understand and respond to customer queries effectively, adapting its communication style based on user feedback.
  • Task Allocation: Assigning tasks appropriately based on the strengths and weaknesses of both humans and AI is critical. Dynamic task allocation, where responsibilities shift based on real-time needs, is often a hallmark of successful HAIC [3]. A good example is in autonomous vehicles, where AI systems may be responsible for navigation under normal conditions, but the human driver takes control when faced with unexpected situations.
@@ 2.3 Modes of Human-AI Collaboration
In the Human-centric mode, humans retain the primary decision-making authority, utilizing AI as an augmentative tool to enhance human capabilities without superseding the human role. This mode values human intuition and oversight, with AI employed to manage data-intensive or repetitive tasks. It optimizes the unique cognitive abilities of humans for strategic and ethical decision-making while leveraging AI to augment human capabilities.
@@
A key example of the symbiotic mode is the Learning to Defer paradigm. This paradigm designs AI systems to understand their own limitations, knowledge, and expertise. Whether confronted with ambiguity or circumstances in which their confidence in a decision is low, the AI system proactively defers to human judgment  [36]. Several studies have explored this concept in depth. Mozannar and Sontag (2020) [37] proposed a Bayesian framework that emphasizes the integration of human and machine predictions. It is crucial to consider potential biases in both agents. This approach aligns with a collaborative decision-making approach because it strives for optimization. It is possible to make decisions by leveraging the strengths of both humans and AI. In a similar way, Mao et al. (2023) [38] extended this concept to scenarios with multiple experts, highlighting the flexibility required in real-world HAIC systems, where multiple sources of expertise may be available. Their work underscores the adaptability of the HAIC systems’ participants, showcasing the ability to dynamically modify task allocation according to the availability and expertise of various agents.
@@
In this section, we looked at the fundamentals of Human-AI Collaboration, such as its main components and the various modes of operation. Evaluating HAIC is critical to ensuring that these systems are successful, efficient, and ethical.
@@
3 Evaluation of Human-AI Collaboration in the Research Literature
@@
{adjustwidth}
-1cm-1cm Reference Domain AI Technology/Method Qn Ql Mix Additional Info Vollmuth et al. (2023) [41] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score. Focuses on human specialists to evaluate AI and HAIC performance. Van Leeuwen et al. (2022) [42] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics. Timmons et al. (2023) [43] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies. Farić et al. (2024) [44] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists. Calisto et al. (2022) [45] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods. Rezwana et al. (2023) [46] Design Generative AI Develops interaction design framework using user feedback. Tschandl et al. (2020) [10] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments. Siu et al. (2021) [47] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed. Reddy et al. (2021) [48] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives. Sharma et al. (2023) [49] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback. Yik et al. (2024) [50] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context. Rastogi et al. (2023) [51] General (Auditing) LLM Examines auditing collaboration using qualitative assessments. Zhao et al. (2023) [52] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics. Makarius et al. (2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods. Aleid et al. (2023) [53] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures. Al-Fatlawi et al. (2023) [54] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives). Yang et al. (2022) [55] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction. Aher et al. (2023) [56] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics. Verheijden & Funk [57] (2023) Design Generative AI Enhances design processes through qualitative co-creation feedback. Hauptman et al. (2023) [58] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments. Sankaran et al. (2022) [59] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics. Massaro (2022) [60] Manufacturing (Safety) Artificial Neural Networks Evaluates safety and risk through quantitative measures. Oravec (2023) [61] Education Large Language Models Discusses academic integrity using qualitative methods. El-Zanfaly et al. (2022) [62] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback. Arias-Rosales (2022) [63] Design/Art Shape Generation Combines quantitative shape metrics with qualitative value assessments. Dziorny et al. (2022) [64] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback. Vössing et al. (2022) [65] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics. Chakravorti et al. (2022) [66] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets. Huang & Rust (2022) [67] Marketing Machine Learning Develops collaborative AI framework using mixed methods. Sachan et al. (2024) [68] Finance Evidential Reasoning Reduces decision noise through quantitative measures. Basu et al. (2021) [69] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data. Dikmen & Burns (2022) [70] Finance (Peer-to-Peer Lending) Explainable AI Examines trust and performance using mixed domain knowledge assessments. Ma et al. (2024) [71] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics. Alon-Barkat & Busuioc (2023) [72] Public Sector Regression Models Explores public sector interactions through mixed methods. Nikolopoulou (2024) [73] Education Generative AI Assesses generative AI’s impact on pedagogy using mixed methods.
Table 1: Generalized Evaluation of Human-AI Collaboration Across Sectors
3.1 Advancements in Evaluation Frameworks
Human-AI Collaboration is being judged by a more traditional metrics. These include ethical concerns, the ability to scale and the ability to adapt different domains.
Dellermann et al. (2021) [25] propose a classification system for hybrid intelligence, emphasizing the importance of understanding the nature of tasks and the dynamic relationship between humans and AI. To categorize different types of HAIC systems, their framework takes into account factors such as level of autonomy, task interdependence, and degree of human control. This classification scheme serves as a foundation for developing tailored evaluation approaches that address the unique characteristics of each type of collaboration.

### v2

@@ Abstract
The use of artificial intelligence (AI) in collaborative working environments, termed Human-AI Collaboration (HAIC), has become vital across diverse domains, enhancing decision-making, efficiency, creativity, and innovation. Despite its wide potential, evaluating HAIC effectiveness poses significant challenges due to the complex, dynamic, and reciprocal interactions between humans and AI systems. This paper provides a comprehensive analysis of existing HAIC evaluation methods and introduces a novel, domain-agnostic framework for systematically assessing these systems. To support structured assessment, our framework incorporates a decision-tree-based approach that helps select relevant metrics tailored to distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic). By integrating both quantitative and qualitative evaluation criteria, the framework aims to address key gaps in HAIC assessment methodologies. We highlight the unique challenges posed by evaluating creative and linguistic AI applications, such as large language models and generative AI in the arts, underscoring the need for tailored evaluation approaches in these emerging areas. This work lays the foundation for future research and empirical validation, offering a structured methodology to enhance the evaluation of HAIC systems.
Keywords Collaboration  $\cdot$ Evaluation  $\cdot$ Symbiosis  $\cdot$ Metrics  $\cdot$ Artificial Intelligence  $\cdot$ Interaction  $\cdot$ Adaptability
@@ 1 Introduction
Evaluating Human-AI Collaboration (HAIC) remains a significant challenge due to the complex interactions between humans and AI systems. The effectiveness of HAIC depends not only on system performance but also on the quality of human-AI interaction, trust, and adaptability. However, existing evaluation methods often fail to capture these multidimensional aspects, making it difficult to establish comprehensive assessment criteria.
To address this issue, a structured framework that provides a foundation for systematically assessing HAIC across various domains, is introduced in the paper. Is is inspired by existing HAIC evaluation methodologies, while it aims at integrating and complement them. While the framework has not yet been applied in real-world scenarios, it is designed to be adaptable and to serve as a reference for future empirical validation.
Section 2 presents a literature review on HAIC evaluation methods, summarizing different approaches, key metrics, and domain-specific challenges. This review highlights gaps in existing evaluation strategies, particularly their inability to fully capture the dynamic and reciprocal nature of HAIC interactions. Based on the finding of this review and the shortcoming identified we introduce an structured evaluation framework [bolman2013reframing] that systematically assesses HAIC across multiple domains (Sections 3– 4) . This framework integrates key evaluation factors such as goals, interaction quality, and task allocation, ensuring a standardized and adaptable assessment tool for practitioners and researchers.
1.1 Background
The rapid advancement of Artificial Intelligence (AI) has transformed multiple domains, leading to a shift from traditional Human-Machine Interaction (HMI) toward Human-AI Collaboration (HAIC) [1]. Unlike conventional HMI, which emphasizes usability and task performance [2, 3], HAIC involves a symbiotic partnership where humans and AI work together to achieve shared goals.
This paradigm shift has demonstrated immense potential across diverse sectors, including manufacturing [4, 5], financial services [6, 7], healthcare [8, 9], and education [10]. However, despite its widespread adoption, evaluating HAIC remains an unresolved challenge. Traditional assessment frameworks, which focus on efficiency and accuracy, are insufficient to capture the complexity of human-AI interaction, adaptability, and trust.
This limitation motivates the need for a comprehensive review of HAIC evaluation methodologies (Section 2) and the development of a structured framework that accounts for these multidimensional aspects (Sections 3– 4).
1.2 Challenges
While HAIC offers numerous opportunities, evaluating its effectiveness remains a complex and unresolved challenge. Recent work by Woelfle et al. (2024) [11] highlights the limitations of traditional evaluation metrics such as task performance, response time, and user satisfaction. Their study emphasizes the need for more comprehensive benchmarks that account for the dynamic and reciprocal nature of HAIC interactions, incorporating metrics like evidence appraisal quality, trust, and adaptability. Such advancements aim to better reflect the multifaceted and collaborative dimensions of human-AI partnerships.
One of the primary challenges lies in assessing the individual and joint contributions of humans and AI systems to collaborative outcomes. Unlike traditional human-machine paradigms, HAIC involves a deeper integration of human intuition and AI capabilities. Evaluating this synergy requires novel metrics that effectively capture the mutual influences between collaborators.
The unique challenges posed by creative and linguistic domains further highlight the limitations of traditional evaluation methods. Applications such as Large Language Models (LLMs) and generative AI in the arts demonstrate the transformative potential of AI to enhance creativity and human expression. However, evaluating these systems requires specialized metrics that move beyond task-based measures. For instance, assessing the interplay of human and AI contributions in artistic creation or the ethical implications of AI-generated content demands methodologies that capture qualitative and contextual nuances. While this study provides a foundational framework for HAIC evaluation, exploring the specific requirements of creative and linguistic domains is an essential avenue for future research.
Additionally, measuring the quality of human-AI interactions remains a critical yet underexplored area. Interaction quality encompasses shared decision-making processes, communication dynamics, and mutual adaptability. Evaluating the ability of both the AI system and human participants to respond to evolving circumstances is crucial for understanding the resilience and effectiveness of HAIC systems in dynamic and high-stakes environments.
1.3 Proposed Approach
Drawing inspiration from established frameworks [12, 2], our approach emphasizes ethical considerations, adaptability, and domain-specific customization. The proposed framework provides a comprehensive structure for assessing HAIC across various modes of interaction, from AI-centric to symbiotic collaborations, integrating quantitative and qualitative measures. Furthermore, it aims to assist researchers and practitioners in identifying existing gaps and opportunities for improving collaboration between humans and AI. By incorporating domain-specific insights, the framework is designed to address practical challenges and applications across fields such as healthcare, finance, and education, thereby demonstrating its versatility and relevance in diverse contexts.
1.4 Contributions
This paper makes the following key contributions:
  • A review of existing HAIC evaluation methodologies, identifying key gaps and challenges that limit their applicability across diverse domains.
  • The introducion a structured evaluation framework for HAIC, integrating both quantitative and qualitative metrics to assess performance, interaction quality, and task allocation.
  • The framework is domain-agnostic, designed to support AI-centric, human-centric, and symbiotic collaboration models, offering a flexible structure applicable to various application areas.
2 Literature Review
This section first examines the transition from HMI to HAIC, highlighting key evaluation challenges. It then explores the fundamental elements that define HAIC interactions, providing the foundation for developing a structured evaluation framework.
@@ 2.1 From Human-Machine Interaction to Human-AI Collaboration
The rapid advancement of HAIC necessitates a broader evaluation perspective that goes beyond traditional HMI usability assessments [13, 14, 15]. HMI methods have primarily been designed to measure user experience and task performance, relying on metrics such as response time, accuracy, and efficiency. However, HAIC introduces new challenges that extend beyond interface usability, requiring metrics that capture interaction quality, collaboration effectiveness, and trust.
Okamura and Yamada (2020) [16] emphasize the importance of system transparency in maintaining trust calibration between humans and AI agents. Over-trusting AI can lead to automation bias, while insufficient trust may result in underutilization. Similarly, Yue et al. (2023) [17] examine how different human-AI collaboration models influence user perceptions and adoption, reinforcing the need for evaluation methods that incorporate trust, interpretability, and adaptability.
2.2 Core Elements and Modes of HAIC
Human-AI Collaboration (HAIC) is defined by the cooperative partnership between individuals and AI systems to accomplish shared objectives. Understanding its key elements is essential for designing effective evaluation methodologies.
@@ 2.2 Core Elements and Modes of HAIC
  • Tasks: HAIC systems support diverse tasks, ranging from automated decision-making to creative problem-solving. The nature of a given task dictates the required level of collaboration between humans and AI [18, 19]. For instance, in healthcare, an AI system may analyze medical images, while a radiologist interprets and integrates results into a final diagnosis.
  • Goals: Collaboration in HAIC is driven by individual and collective goals. AI objectives often include enhancing efficiency and accuracy, while human goals may involve skill augmentation and contextual reasoning [20]. Continuing with the radiology example, the AI assists in accurate disease detection, while radiologists focus on clinical decision-making and patient-centered care.
  • Interaction: The effectiveness of HAIC is determined by communication quality and feedback mechanisms [21, 22]. Misalignment in expectations or miscommunication can lead to ineffective collaboration. For instance, AI-driven customer support chatbots must interpret user queries and refine responses dynamically to ensure effective problem resolution.
  • Task Allocation: Dynamic task allocation is essential in HAIC, ensuring that responsibilities shift based on real-time conditions and expertise distribution [23]. For example, in advanced driver assistance systems, AI controls navigation under normal conditions, but the human driver intervenes when unexpected obstacles arise.
@@
In the Human-centric mode, humans retain the primary decision-making authority, utilizing AI as an augmentative tool to enhance human capabilities without superseding the human role. This mode values human intuition and oversight, with artificial intelligence employed to manage repetitive or data-intensive tasks. It optimizes the unique cognitive abilities of humans for strategic and ethical decision-making while leveraging AI to augment human capabilities.
@@
A key example of the symbiotic mode is the Learning to Defer (L2D) paradigm, where AI systems are designed to recognize their own limitations and defer decisions to human judgment when necessary. This approach enhances collaborative decision-making by leveraging the strengths of both humans and AI. Recent studies have expanded on this concept. Alves et al. (2024) [31] introduced a cost-sensitive L2D framework that considers scenarios with multiple experts and workload constraints, emphasizing the need for flexibility in real-world Human-AI Collaboration (HAIC) systems. Additionally, Hemmer et al. (2022) [32] proposed an approach that trains a classification model to complement the capabilities of multiple human experts. Their work underscores the adaptability of the HAIC systems’ participants, showcasing the ability to dynamically modify task allocation according to the availability and expertise of various agents.
@@
In this section, we looked at the fundamentals of Human-AI Collaboration, such as its main components and the various modes of operation. The evaluation of HAIC is critical to ensure that these systems are successful, efficient, and ethical.
@@
2.3 Evaluation Approaches in HAIC
@@
Table 1: Examples of Evaluation of Human-AI Collaboration Across Sectors
Reference Domain AI Technology/Method Qn Ql Mix Additional Info
Vollmuth et al. (2023) [35] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score. Focuses on human specialists to evaluate AI and HAIC performance.
Van Leeuwen et al. (2022) [36] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics.
Timmons et al. (2023) [37] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies.
Farić et al. (2024) [38] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists.
Calisto et al. (2022) [39] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods.
Rezwana et al. (2023) [40] Design Generative AI Develops interaction design framework using user feedback.
Tschandl et al. (2020) [8] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments.
Siu et al. (2021) [41] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed.
Reddy et al. (2021) [42] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives.
Sharma et al. (2023) [43] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback.
Yik et al. (2024) [44] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context.
Rastogi et al. (2023) [45] General (Auditing) LLM Examines auditing collaboration using qualitative assessments.
Zhao et al. (2023) [46] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics.
Makarius et al. (2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods.
Aleid et al. (2023) [47] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures.
Al-Fatlawi et al. (2023) [48] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives).
Yang et al. (2022) [49] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction.
Aher et al. (2023) [50] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics.
Verheijden & Funk [51] (2023) Design Generative AI Enhances design processes through qualitative co-creation feedback.
Hauptman et al. (2023) [52] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments.
Sankaran et al. (2022) [53] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics.
Massaro (2022) [54] Manufacturing (Safety) Artificial Neural Networks Evaluates safety and risk through quantitative measures.
Oravec (2023) [55] Education Large Language Models Discusses academic integrity using qualitative methods.
El-Zanfaly et al. (2022) [56] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback.
Arias-Rosales (2022) [57] Design/Art Shape Generation Combines quantitative shape metrics with qualitative value assessments.
Dziorny et al. (2022) [58] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback.
Vössing et al. (2022) [59] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics.
Chakravorti et al. (2022) [60] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets.
Huang & Rust (2022) [61] Marketing Machine Learning Develops collaborative AI framework using mixed methods.
Sachan et al. (2024) [62] Finance Evidential Reasoning Reduces decision noise through quantitative measures.
Basu et al. (2021) [63] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data.
Dikmen & Burns (2022) [64] Finance (Peer-to-Peer Lending) Explainable AI Examines trust and performance using mixed domain knowledge assessments.
Ma et al. (2024) [65] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics.
Alon-Barkat & Busuioc (2023) [66] Public Sector Regression Models Explores public sector interactions through mixed methods.
Nikolopoulou (2024) [67] Education Generative AI Assesses generative AI’s impact on pedagogy using mixed methods.
2.3.1 Advancements in Tools, Methodologies, and Real-Time Collaboration for HAIC
Human-AI Collaboration (HAIC) is increasingly evaluated using both traditional and innovative metrics, with an emphasis on ethical concerns, scalability, and adaptability across domains. Recent advancements in tools and methodologies are reshaping the evaluation landscape, offering novel ways to understand and optimize human-AI interactions.
Fischer (2023) [68] envisions the future of human-AI collaboration as a deeply intertwined partnership, where AI systems are designed not only to assist with tasks but also to foster positive emotional and social interactions. His work explores scenarios where AI colleagues are capable of understanding and responding to human emotions, aiming to enhance collaboration by creating a sense of trust and mutual satisfaction. This forward-looking perspective emphasizes the need for evaluation frameworks that address not only technical performance but also the social and emotional dimensions of human-AI interactions. Fischer’s work highlights the importance of incorporating human-centered design principles into AI systems to promote effective collaboration and user well-being.